OpenStudy (anonymous):

Can someone please teach me step by step. Solve the equation for w. 3(w + 7) – 5w = w + 12

OpenStudy (anonymous):

Sure!

OpenStudy (anonymous):

First open the brackets on 3(x+7).

OpenStudy (anonymous):

Once you do that, put them together with -5w.

OpenStudy (anonymous):

so 3w+7-5w

OpenStudy (anonymous):

Right!

OpenStudy (anonymous):

So you have -2w+7=w+12

OpenStudy (anonymous):

w=3

OpenStudy (anonymous):

@Tusharsingh Don't give direct answers!! It's against the Code of Conduct

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I already knew the answer but I just don't know how to get it.

OpenStudy (anonymous):

Thats why I wanted step by step lol

OpenStudy (anonymous):

Just put the w on one side of the equation and the numbers on the other side of the equation.

OpenStudy (anonymous):

Can you write it out for me please.

OpenStudy (anonymous):

-2w+12w=7

OpenStudy (anonymous):

Actualy you have -2w+21=w+12 becuase of 3(x+7)=3x+21-5w -3w=-9 w=3

OpenStudy (anonymous):

So, do you get it or you're still stuck?

OpenStudy (anonymous):

I'm still stuck.

OpenStudy (anonymous):

Do you know of a website that would be able to help me with these problems?

OpenStudy (anonymous):

Yes https://mathway.com/

OpenStudy (anonymous):

You can sign up of a membership and they will give you step-by-step answers. I use it a lot to confirm my answers are correct.

OpenStudy (anonymous):

Okay Thank you!

OpenStudy (anonymous):

Your welcome!